# TrustGeoGen: Formal-Verified Data Engine for Trustworthy Multi-modal Geometric Problem Solving

### v2

@@ Abstract
Mathematical geometric problem solving (GPS) demands verifiable logical coherence and multimodal reasoning capabilities. While large language models (LLMs) have shown rapid progress in GPS, their advancement is hindered by the lack of reliable benchmarks and systematic methodologies. A critical challenge is the inherent hallucination in LLMs, which leads to synthetic GPS datasets that are often noisy, unverified, and self-contradictory. To address this, we introduce TrustGeoGen, a data engine that generates formally verified geometric problems to establish a principled and trustworthy benchmark. Our engine integrates four key innovations: 1) Multimodal Alignment, which synchronizes the generation of diagrams, text, and step-by-step solutions; 2) Formal Verification, ensuring all reasoning paths are rule-compliant; 3) Connection Thinking, bridging formal deduction with human-like logical steps; and 4) our GeoExplore series algorithms, which produce diverse problem variants with multiple solutions and self-reflective backtracking. Using this engine, we create the GeoTrust-200K dataset and the corresponding GeoTrust-test benchmark, both with guaranteed cross-modal integrity. Experiments reveal that state-of-the-art models achieve only 45.83% accuracy on GeoTrust-test, highlighting its significant challenge. Furthermore, training on our synthesized data substantially improves model performance on GPS tasks, with strong generalization to out-of-domain (OOD) benchmarks. Our code and data are available at https://github.com/Alpha-Innovator/TrustGeoGen.
Index Terms:
Data Engine, Geometry, Multi-modal
1 Introduction
Geometric problem solving (GPS) geox/Xia2024GeoXGP , as a pivotal branch of mathematical reasoning or more broadly speaking, general reasoning, demands tripartite capabilities: 1) extracting spatial relationships from visual diagrams, 2) performing mathematical reasoning in language modality, and 3) maintaining stepwise logical coherence where intermediate conclusions rigorously justify subsequent deductions. The development of trustworthy AGI systems fundamentally relies on reliably annotated multimodal data that not only bridges visual-textual modalities but also preserves closed-loop reasoning chains. This requirement stems from a critical axiom: a geometrically valid solution demands both correct answers and air-tight logical pathways, as human experts reject proofs with leap-of-faith [^1] Here it means an unjustified or unsubstantiated jump in reasoning, where a conclusion is reached with insufficient logical steps or evidence. arguments even when final results are accurate.
The recent advancement of multimodal large language models (MLLMs) (math-llava/shi2024math, ; mathvista/lumathvista, ; peng2024chimera, ; qwen-vl/bai2023qwen, ; openai2024gpt4o, ; g-llava/gao2023g, ; chen2024far, ; lu2024deepseek, ; geox/Xia2024GeoXGP, ; wu2024deepseekvl2mixtureofexpertsvisionlanguagemodels, ; internlmxcomposer2, ; internlmxcomposer2_4khd, ) has revealed promising capabilities in geometric reasoning, demonstrating competence in solving elementary geometric problems. Parallel developments in external formalized expert systems (unigeo_geoformer/chen2022unigeo, ; unimath/liang2023unimath, ; pgpsnet/zhang2023multi, ; inter-gps/lu2021inter, ) have extended problem solving capabilities to more complex geometric configurations, with highly specialized models (alphageometry/Trinh2024, ; alphageometry2/chervonyi2025, ; tonggeometry/Zhang2024ProposingAS, ) even approaching challenges of the international mathematical olympiad (IMO) level. However, these achievements are critically dependent on high-quality geometric data, currently in short supply. The existing datasets are mainly derived from manual annotations of secondary school textbook problems (geoqa_ngs/chen2021geoqa, ; unigeo_geoformer/chen2022unigeo, ; pgdp5k/hao2022pgdp5k, ; pgpsnet/zhang2023multi, ; inter-gps/lu2021inter, ), and this process is limited in scale and diversity. Although recent integration of MLLMs (e.g. GPT series (gpt4/achiam2023gpt, ; openai2023gpt4v, ; openai2024gpt4o, )) for synthetic question-answer generation (mavis/zhang2024mavis, ; g-llava/gao2023g, ) has partially ameliorated data scarcity concerns, significant challenges persist in ensuring stepwise solution validity [^2] Specifically, automatically generated procedural labels frequently contain latent inconsistencies, such as misapplied geometric axioms or diagram-coordinate mismatches., which may undermine the logical coherence of reasoning chains during model training.
@@ 1 Introduction
GeoQA geoqa_ngs/chen2021geoqa 4998 I, NS, A Middle School D Human Annotation
GeoQA+ dpe-gps/cao2022augmented 7528 I, NS, A Middle School D Human Annotation
Geometry3K inter-gps/lu2021inter 3002 I, FS, A Middle School D Step-wise Formal Verified
PGPS9K pgpsnet/zhang2023multi 9022 I, FS, A Middle School D Step-wise Formal Verified
UniGeo unigeo_geoformer/chen2022unigeo 9543 I, NS, A Middle School D From Source File
GeoEval zhang2024geoeval 5050 I, A Middle & High School - LLM Labeling
MathVista mathvista/lumathvista 1320 * I, A Middle School - From Source File
MathVerse zhang2024mathverse 1745 * I, NS, A High School - LLM Labeling
GeoSense leonardi2024geosence 1789 I, NS, A Synthetic D LLM Labeling & Human Annotation
OlympiadBench he2024olympiadbench 8476 I, NS, A Olympiad-level D From Source File
MAVIS mavis/zhang2024mavis 588K I, NS, A Synthetic D LLM Labeling
G-LLAVA g-llava/gao2023g 170K I, NS, A Synthetic D LLM Lableling
@@
GeoTrust-test 240 I, FS, NS, A Synthetic D Step-wise Formal Verified
@@
However, as depicted in table˜I, current geometric datasets exhibit four deficiencies: 1) Modality Fragmentation: The existing datasets suffer from incomplete modality alignment, lacking annotations that synchronize images, textual captions, questions, and chain-of-thought reasoning processes within unified instances. 2) Scalability Issues: Most approaches rely on repurposing human-authored problems from textbooks or examinations rather than generating novel, curriculum-aligned challenges. 3) Cost Constraints: Manual annotation requires costly geometry expertise, while generating solutions through MLLMs consumes excessive computational resources. 4) Trust Deficits: Model-generated solutions lack reliable automated verification mechanisms, introducing risks of undetected logical errors or misaligned reasoning steps.
To this end, we develop TrustGeoGen, a formal language-verified geometric data generation engine that autonomously produces geometric problems, diagrams, and stepwise solutions through four integrated components: 1) a Constructor to generate constraint-satisfying problem premises and diagrams, 2) a Reasoner expanding geometrically valid reasoning graphs with rigorous node verification, 3) a Sampler leveraging the algorithm GeoExplore to extract high-quality reasoning paths, which are later converted into problems by pairing initial premises with derived conclusions, and 4) A Translator converting the data constructed by the formal engine into natural language, and establishing "connection thinking" to bridge the gap between formal reasoning steps and natural human logic. Furthermore, we improve the problem complexity through a bootstrapping mechanism in TrustGeoGen, where formally verified geometric premises in prior stages are recursively fed back to the Constructor to iteratively append new premises. This closed-loop process enables systematic escalation of difficulties while maintaining logical consistency. Additionally, we refine the algorithm GeoExplore to simultaneously explore alternative solution paths and self-correcting reasoning trajectories, generating problems with multiple valid solutions and self-reflective traceback data.
Our proposed TrustGeoGen ensures: 1) Modality Completeness with synchronized dual-form (formal & informal) captions, questions, solutions, and diagrams, 2) Scalability in generating geometry problems by bootstrapping mechanisms without reliance on existing data source and 3) Verifiability through formal validation of all solution steps in its integrated reasoner. Leveraging TrustGeoGen, we synthesized a raw dataset of 200K problems, from which we curated the GeoTrust-train set (2,342 samples) and the GeoTrust-test set (240 samples).
Figure 1: TrustGeoGen overview: (1) Constructor builds geometric premises from a base scene by incrementally adding clauses and enforces constraints via a compiler, (2) Reasoner generates rigorously validated reasoning graph from premises by geometric rules, (3) Sampler selects conclusion nodes from the graph and traces reasoning paths to formalize problems and solutions, (4) Translator converts formal language to natural language for metadata. A bootstrap mechanism further enhances data complexity and problem difficulty.
Our experimental results demonstrate that the data generated by TrustGeoGen not only reaches a difficulty level comparable to the international mathematical olympiad (IMO) but also contains novel problems unseen in the pre-training corpora of state-of-the-art (SOTA) models, thereby enhancing their performance. Furthermore, the "connection thinking" provided by the translator bridges the gap between formal engine-verified reasoning paths and human-like logical inference, significantly boosting the model’s geometric problem-solving abilities. The data constructed from "thinking templates" by the GeoExplore series of algorithms has also been proven effective in improving the model’s reasoning capabilities. Additionally, our training data from GeoTrust enhances model performance on the out-of-distribution (OOD) testset GeoQA (geoqa_ngs/chen2021geoqa, ), Geometry3K inter-gps/lu2021inter , and OlympiadBench he2024olympiadbench , highlighting its broader generalization capability. The main contributions of this paper can be summarized as follow:
@@
  • We develop TrustGeoGen, a formalized geometric data generation engine, producing multimodal-integrated data (geometric diagrams, formal & natural language captions, questions and solutions) with trustworthy reasoning guarantees via formal verification.
  • We introduce "Connection Thinking" to bridge the gap between formal and human-like logical reasoning, enabling our engine to enhance models’ reasoning capabilities. Our GeoExplore series algorithms further synthesize data with diverse thinking templates, offering different problem-solving pathways.
  • Extensive experiments show that TrustGeoGen generates challenging, verifiable data that significantly improves models’ geometric problem-solving abilities. Results on OOD datasets demonstrate the effectiveness and generalization of our synthetic data.
2 Related Work
2.1 Advances in Multimodal Large Language Model
@@
Recent advances in large language models (LLMs) have demonstrated remarkable progress in linguistic intelligence, achieving human-level capabilities in various applications (chatgpt/ouyang2022training, ; llama/touvron2023llama, ; llama2/touvron2023llama, ; internlm/team2023internlm, ). Building upon this foundation, the research community has focused on extending these text-based architectures to process visual information, leading to the emergence of sophisticated multimodal large language models (MLLMs) (qwen-vl/bai2023qwen, ; gpt-4v/achiam2023gpt, ; reid2024gemini, ). These integrated systems typically employ modality alignment mechanisms, such as Q-former (blip-2/li2023blip, ) and linear layers (llava/liu2024visual, ), to establish connections between visual representations and text embeddings. Recently, an important application of MLLMs is their ability to be regarded as multimodal agents that interact with scientific environments (team2025novelseek, ; zhang2025origene, ; gottweis2025towards, ), thereby generating valuable scientific discoveries and driving productivity advancements. Although current MLLMs show strong performance in standard vision-language benchmarks (fu2023mme, ; xia2024docgenome, ; structchart/xia2023structchart, ; mathvista/lumathvista, ; jiang2025mme, ), their effectiveness decreases significantly when processing mathematical visualizations that require reasoning. Recent specialized approaches address this limitation through targeted training strategies. For example, geometric reasoning capabilities have been enhanced through domain-specific datasets containing annotated diagrams (g-llava/gao2023g, ; mavis/zhang2024mavis, ).
2.2 Geometric Problem Solving
@@
As a challenging task, geometric problem solving requires understanding diagrams, interpreting symbols, and performing complex reasoning. While MLLMs have shown remarkable proficiency in human-level reasoning and generation capabilities, they struggle with automatic geometric problem solving due to their pre-training on natural images and texts, as well as the lack of automated verification in the problem-solving process. To address this limitation, researchers have proposed various approaches to improve the understanding of geometric images and corpora, including unimodal pre-training, vision-language alignment, and visual instruction tuning (geox/Xia2024GeoXGP, ; mavis/zhang2024mavis, ; g-llava/gao2023g, ; geoqa_ngs/chen2021geoqa, ; unigeo_geoformer/chen2022unigeo, ; pgpsnet/zhang2023multi, ; pgdp5k/hao2022pgdp5k, ; jiang2025mme, ). Another major approach involves the use of formalized solvers to tackle geometric problems. While these solvers (alphageometry/Trinh2024, ; alphageometry2/chervonyi2025, ; sicca2024newclid, ) are capable of addressing challenges at the level of the international mathematical olympiad (IMO), they require precise modeling of geometric problems to fit the solver’s language. This necessity introduces significant obstacles in terms of practical application and generalization, limiting the usability in real-world scenarios.
2.3 Datasets and Benchmarks
@@
High-quality training data is crucial for enhancing the geometry problem-solving capabilities of MLLMs. As summarized in table˜I, there are three common data construction approaches. The first approach involves filtering real-world data and then annotating it manually geoqa_ngs/chen2021geoqa ; dpe-gps/cao2022augmented ; unigeo_geoformer/chen2022unigeo ; mathvista/lumathvista ; he2024olympiadbench . While this method yields high-quality multi-modal reasoning data, its scalability is limited by the finite data pool and the intensive labor required. Furthermore, as problem complexity increases, the difficulty of annotation rises sharply, typically constraining the resulting problems to a lower difficulty range. The second approach utilizes formal engines for data generation inter-gps/lu2021inter ; pgpsnet/zhang2023multi . This allows for the rapid creation of large-scale datasets with verifiably correct reasoning processes. However, the quality of the data is inherently tied to the sophistication of the formal engine. A significant gap often exists between the synthesized logic and authentic human reasoning, leaving considerable room for improvement in enhancing the model’s natural language inference abilities. Finally, the third method employs LLMs to synthesize reasoning data mavis/zhang2024mavis ; g-llava/gao2023g . This method is highly efficient, and the generated reasoning steps more closely resemble natural language. Its primary drawback, however, is that the outputs are unverifiable; they may contain erroneous intermediate steps that could mislead or corrupt the model’s reasoning capabilities.
3 Methodology

### v3

@@ Abstract
Geometric problem solving (GPS) requires precise multimodal understanding and rigorous, step-by-step logical reasoning. However, developing capable Multimodal Large Language Models (MLLMs) for GPS is heavily bottlenecked by the scarcity of high-quality, verifiable data. Existing data acquisition paradigms either suffer from modality incompleteness and unverified logical gaps ("leaps-of-faith"), or rely on formal engines that generate rigid, structurally homogeneous data, failing to produce high-difficulty problems or foster genuine natural-language reasoning. To overcome these limitations, we introduce TrustGeoGen, an autonomous and formalized geometric data generation engine. TrustGeoGen strictly guarantees reasoning trustworthiness through formal verification while generating multimodally integrated data, including premises, visual diagrams, and solutions. To systematically scale problem difficulty, we incorporates difficulty-aware filtering and iterative bootstrapping mechanism. Furthermore, we propose "connection thinking" to bridge the semantic gap between rigid formal logic and fluent human-like reasoning, ensuring coherent logical transitions. We also introduce the GeoExplore family of sampling algorithms to extract diverse problem-solving trajectories based on various thinking templates. Extensive experiments demonstrate that training models on our synthesized dataset, GeoTrust, substantially enhances deep geometric reasoning capabilities and yields significant performance gains across out-of-distribution (OOD) benchmarks, including GeoQA, Geometry3K, and OlympiadBench.
I Introduction
Geometric problem solving (GPS) [geox/Xia2024GeoXGP, peng-etal-2023-geodrl, wu2024gps] is a representative task for mathematical reasoning. It requires models to jointly interpret diagrams, understand textual conditions, and produce logically valid multi-step solutions, where each conclusion should be justified by prior statements [feng2025geobench]. As a result, GPS is a challenging benchmark for trustworthy multimodal reasoning: a correct final answer alone is insufficient if the intermediate reasoning is incomplete or unjustified. Progress on this task therefore depends on multimodal data that aligns diagrams, language, and verifiable reasoning steps.
Recent multimodal large language models (MLLMs) [math-llava/shi2024math, mathvista/lumathvista, peng2024chimera, qwen-vl/bai2023qwen, openai2024gpt4o, g-llava/gao2023g, chen2024far, lu2024deepseek, geox/Xia2024GeoXGP, wu2024deepseekvl2mixtureofexpertsvisionlanguagemodels, internlmxcomposer2, internlmxcomposer2_4khd] have shown encouraging progress on elementary geometry, while formalized expert systems [unigeo_geoformer/chen2022unigeo, unimath/liang2023unimath, pgpsnet/zhang2023multi, inter-gps/lu2021inter] and specialized solvers [alphageometry/Trinh2024, alphageometry2/chervonyi2025, tonggeometry/Zhang2024ProposingAS] have demonstrated strong potential on more challenging problems. However, existing pipelines still do not provide models with sufficiently coherent multimodal reasoning supervision. In particular, when language models are used mainly to interact with formal solvers, the resulting data may be formally correct but linguistically rigid, making it difficult for models to learn natural, well-connected reasoning explanations. This limitation highlights the need for training data that is both formally verifiable and natural-language coherent.
As shown in Table I, current multimodal geometry data mainly comes from two paradigms: annotating existing problems [geoqa_ngs/chen2021geoqa, dpe-gps/cao2022augmented, zhang2024geoeval, leonardi2024geosence] and synthesizing data with formal engines [inter-gps/lu2021inter, pgpsnet/zhang2023multi]. Annotation-based approaches rely on textbook problems or curated datasets, with solutions written manually or assisted by LLMs. While practical, they are constrained by the source data and often suffer from incomplete modalities or missing intermediate steps, especially for difficult problems. Formal synthesis, in contrast, can produce stepwise verifiable reasoning at much lower annotation cost. Yet existing synthetic pipelines typically lack systematic difficulty control, and the generated explanations are often template-like, weakly connected across steps, and limited in reasoning diversity. Consequently, existing datasets still leave substantial room for improvement in completeness, verifiability, and reasoning quality.
@@ I Introduction
GeoQA [geoqa_ngs/chen2021geoqa] 4998 I, NS, A Middle School D Human Annotation
GeoQA+ [dpe-gps/cao2022augmented] 7528 I, NS, A Middle School D Human Annotation
Geometry3K [inter-gps/lu2021inter] 3002 I, FS, A Middle School D Step-wise Formal Verified
PGPS9K [pgpsnet/zhang2023multi] 9022 I, FS, A Middle School D Step-wise Formal Verified
UniGeo [unigeo_geoformer/chen2022unigeo] 9543 I, NS, A Middle School D From Source File
GeoEval [zhang2024geoeval] 5050 I, A Middle & High School - LLM Labeling
MathVista [mathvista/lumathvista] 1320 * I, A Middle School - From Source File
MathVerse [zhang2024mathverse] 1745 * I, NS, A High School - LLM Labeling
GeoSense [leonardi2024geosence] 1789 I, NS, A Synthetic D LLM Labeling & Human Annotation
OlympiadBench [he2024olympiadbench] 8476 I, NS, A Olympiad-level D From Source File
MAVIS [mavis/zhang2024mavis] 588K I, NS, A Synthetic D LLM Labeling
G-LLAVA [g-llava/gao2023g] 170K I, NS, A Synthetic D LLM Lableling
AutoGeo [huang2025autogeo] 100K I, FS, NS, A Synthetic D Step-wise Formal Verified
@@
GeoTrust-test 240 I, A Synthetic - Step-wise Formal Verified
@@
To address these issues, we propose TrustGeoGen, a formally verified generation engine for multimodal geometric data. TrustGeoGen consists of four components: (1) a Constructor that generates geometric premises and corresponding diagrams under explicit constraints; (2) a Reasoner that expands formally valid reasoning graphs through rule-based verification; (3) a Sampler that uses GeoExplore to extract high-quality reasoning paths and formulate problem-solution pairs; and (4) a Translator that converts formal derivations into fluent natural-language explanations. To generate more challenging and informative data, TrustGeoGen further incorporates difficulty-aware filtering and iterative bootstrapping. In addition, we introduce connection thinking, which bridges the semantic gap between rigid formal logic and fluent human-like reasoning, ensuring coherent logical transitions, and use the GeoExplore family of sampling algorithms to extract diverse problem-solving trajectories based on various thinking templates.
Experiments show that TrustGeoGen can generate challenging geometry data and that training on the resulting dataset consistently improves model performance. The proposed translation strategy helps narrow the gap between formally verified derivations and human-readable reasoning, while the GeoExplore sampling algorithms further improve reasoning diversity. Models trained on the synthesized dataset, GeoTrust, achieve substantial gains on out-of-distribution benchmarks, including GeoQA [geoqa_ngs/chen2021geoqa], Geometry3K [inter-gps/lu2021inter], and OlympiadBench [he2024olympiadbench], demonstrating strong generalization to unseen geometry problems. Our main contributions are as follows:
@@
  • We propose TrustGeoGen, a generation engine for geometric data that produces integrated multimodal outputs, including diagrams, formal descriptions, natural-language explanations, questions, and solutions, with formally verifiable intermediate reasoning steps.
  • We introduce connection thinking to improve the coherence between formal derivations and natural-language explanations, and propose the GeoExplore sampling family to increase the diversity of generated reasoning paths.
  • Extensive experiments show that data generated by TrustGeoGen is challenging, verifiable, and effective for improving geometric problem-solving performance, with strong transfer to out-of-distribution benchmarks.
II Related Work
II-A Advances in Multimodal Large Language Model
@@
Recent advances in large language models (LLMs) have demonstrated remarkable progress in linguistic intelligence, achieving human-level capabilities in various applications [chatgpt/ouyang2022training, llama/touvron2023llama, llama2/touvron2023llama, internlm/team2023internlm]. Building upon this foundation, the research community has focused on extending these text-based architectures to process visual information, leading to the emergence of sophisticated multimodal large language models (MLLMs) [qwen-vl/bai2023qwen, gpt-4v/achiam2023gpt, reid2024gemini]. These integrated systems typically employ modality alignment mechanisms, such as Q-former [blip-2/li2023blip] and linear layers [llava/liu2024visual], to establish connections between visual representations and text embeddings. Recently, an important application of MLLMs is their ability to be regarded as multimodal agents that interact with scientific environments [team2025novelseek, zhang2025origene, gottweis2025towards], thereby generating valuable scientific discoveries and driving productivity advancements. Although current MLLMs show strong performance in standard vision-language benchmarks [fu2023mme, xia2024docgenome, structchart/xia2023structchart, mathvista/lumathvista, jiang2025mme], their effectiveness decreases significantly when processing mathematical visualizations that require reasoning. Recent specialized approaches address this limitation through targeted training strategies. For example, geometric reasoning capabilities have been enhanced through domain-specific datasets containing annotated diagrams [g-llava/gao2023g, mavis/zhang2024mavis].
II-B Geometric Problem Solving
@@
As a challenging task, geometric problem solving requires understanding diagrams, interpreting symbols, and performing complex reasoning. While MLLMs have shown remarkable proficiency in human-level reasoning and generation capabilities, they struggle with automatic geometric problem solving due to their pre-training on natural images and texts, as well as the lack of automated verification in the problem-solving process. To address this limitation, researchers have proposed various approaches to improve the understanding of geometric images and corpora, including unimodal pre-training, vision-language alignment, and visual instruction tuning [geox/Xia2024GeoXGP, mavis/zhang2024mavis, g-llava/gao2023g, geoqa_ngs/chen2021geoqa, unigeo_geoformer/chen2022unigeo, pgpsnet/zhang2023multi, pgdp5k/hao2022pgdp5k, jiang2025mme]. Another major approach involves the use of formalized solvers to tackle geometric problems. While these solvers [alphageometry/Trinh2024, alphageometry2/chervonyi2025, sicca2024newclid] are capable of addressing challenges at the level of the international mathematical olympiad (IMO), they require precise modeling of geometric problems to fit the solver’s language. This necessity introduces significant obstacles in terms of practical application and generalization, limiting the usability in real-world scenarios.
II-C Datasets and Benchmarks
@@
High-quality training data is crucial for enhancing the geometry problem-solving capabilities of MLLMs. As summarized in Table˜I, there are three common data construction approaches. The first approach involves filtering real-world data and then annotating it manually [geoqa_ngs/chen2021geoqa, dpe-gps/cao2022augmented, unigeo_geoformer/chen2022unigeo, mathvista/lumathvista, he2024olympiadbench]. While this method yields high-quality multi-modal reasoning data, its scalability is limited by the finite data pool and the intensive labor required. Furthermore, as problem complexity increases, the difficulty of annotation rises sharply, typically constraining the resulting problems to a lower difficulty range. The second approach utilizes formal engines for data generation [inter-gps/lu2021inter, pgpsnet/zhang2023multi]. This allows for the rapid creation of large-scale datasets with verifiably correct reasoning processes. However, the quality of the data is inherently tied to the sophistication of the formal engine. A significant gap often exists between the synthesized logic and authentic human reasoning, leaving considerable room for improvement in enhancing the model’s natural language inference abilities. Finally, the third method employs LLMs to synthesize reasoning data [mavis/zhang2024mavis, g-llava/gao2023g]. This method is highly efficient, and the generated reasoning steps more closely resemble natural language. Its primary drawback, however, is that the outputs are unverifiable; they may contain erroneous intermediate steps that could mislead or corrupt the model’s reasoning capabilities.
III Methodology